OpenStudy (anonymous):

If the dimensions of a square are increased by 50%, its area is increased by?

OpenStudy (anonymous):

area of a square is given by A=x^2 where x is length of a side of a square now if x is increased by 50%,that means it is increased by (50/100)x=x/2 therefore the new side length will be x+x/2=(3/2)x now find the new area and compare it with the one we had initially